OpenStudy (anonymous):

The equation 0.3x = 7 was translated from the problem "_____." 0.3 is what percent of 7? 30% of what number is 7? What number is 30% of 7? 7 is what percent of 0.03?

OpenStudy (anonymous):

30% of what number is 7?

OpenStudy (anonymous):

Well, 100% of something is "all of it". So 100% of 6 is 1 * 6. 50% of something is half of it. So 50% of 6 is .5 * 6. 25% of something is one quarter of it, so 25% of 6 is 1/4 * 6 = 0.25 * 6. Notice a pattern?

OpenStudy (anonymous):

30% of what number is 7? 30% of x = 7 30/100 * x = 7 0.3x = 7

OpenStudy (anonymous):

30% of some number x is 0.3 * x.